OpenStudy (queenv):

ompounded semiannually. P dollars is invested at annual interest rate r for 1 year. If the interest is compounded semiannually, then the polynomial P(1+r/2)2 represents the value of the investment after 1 year. Rewrite this expression without parentheses. Evaluate the polynomial if P = $5670 and r = 3.5%

OpenStudy (kropot72):

\[P(1+\frac{r}{2})^{2}=P(1+r+\frac{r ^{2}}{4})=P+\Pr+P\frac{r ^{2}}{4}\ .........(1)\] Now plug the values P = 5670 and r = 0.035 into equation (1).

OpenStudy (tkhunny):

And why would anyone EVER do that? We changed it from Addition: 1 Multiplication: 1 Division: 1 Exponentiation: 1 To this Addition: 2 Multiplication: 2 Division: 1 Exponentiation: 1 Can you even imagine how much more complex a larger problem would be under this transformation?

OpenStudy (kropot72):

The question asks "Rewrite this expression without parentheses", presumably as an exercise in expansion.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."